Whats the distance between these two points? T(3,4) and W(2,7)? Is it between 3 and 4??
Answer by stanbon(48558)

Whats the distance between these two points? T(3,4) and W(2,7)? Is it between 3 and 4??
------------------
d^2 = x^2 + y^2
----------------------------------
d^2 = (3-2)^2 + (7-4)^2
---------------------------
d = sqrt[1^2 + 3^2]
---
d = sqrt[10]
====================
Yes, the distance is between 3 and 4
=========================================
Cheers,
Stan H.